how can we solve this in linear equation 5x-3=5+1/2y
Answer by stanbon(75887)

5x-3=5+(1/2)y
5x-8 = (1/2)y
Multiply thru by 2 to get:
y = 10x - 16
===============
Cheers,
Stan H.